OpenStudy (anonymous):

Let R be the region bounded by the x-axis, the graph y=(x+1)^(1/2), and the line x=3. Find the value h such that the vertical line x=h divides the region R into 2 regions of equal area.

OpenStudy (anonymous):

how far have you reached? i will help you further

OpenStudy (tkhunny):

What's your plan. There should be some integrals in your future.

OpenStudy (anonymous):

I have already calculated that the area is 5.3333. So assumed that I should divide that by 2 and work backwards form there?

OpenStudy (tkhunny):

Backwards? No. Solve for a \(\int\limits_{-1}^{a}\sqrt{x+1}\;dx = \dfrac{8}{3}\)

OpenStudy (anonymous):

WHere did the 8/3 come from?

OpenStudy (tkhunny):

The problem statement, "divides the region R into 2 regions of equal area. " The area of the region is 16/3. Half the area is 8/3. Note: I guess I should have used 'h' instead of 'a'. The name is given in the problem statement.

OpenStudy (anonymous):

So would i integrate and place h in the place of x and subtract f(-1)?

OpenStudy (anonymous):

I did this wrong... i got 7

OpenStudy (tkhunny):

Just like you already did when you found the total area. You could also do it the other way, [h,3], instead of [-1,h].

OpenStudy (tkhunny):

Good call. h = 7 clearly is no good.

OpenStudy (anonymous):

Can you slowly go step by step on this. I'm just not use to seeing a letter on the integral

OpenStudy (tkhunny):

How did you get the total area?

OpenStudy (anonymous):

the integral of (x+1)^(1/2) form -1 to 3

OpenStudy (tkhunny):

Please demonstrate this. How did you find the correct value of that integral?